OpenStudy (anonymous):

A partial solution to the equation 8b + 12 = 5(4 + 2b) is given below. Choose the step that is missing from the solution.

OpenStudy (anonymous):

8b + 12 = 5(4 + 2b) 8b = 8 + 10b −2b = 8 b = −4